# Global well-posedness of 2D Navier-Stokes with Dirichlet boundary fractional noise

### v1

@@ 3.2. Auxiliary Navier-Stokes type equations
with $A_{q}$ is the Stokes operator on $\mathbb{L}^{q}$ and $B$ is the bilinear nonlinearity as in Subsection 2.3. Finally, $q=r/2$ and $(w,r)$ satisfies the following
@@
Note that the above assumption is satisfied $\mathbf{P}-a.s.$ with $w=w_{g}$, as it follows from 3.1. Moreover, the limitation $r<4$ is motivated by 3.2. In particular, the arguments used in [1] do not apply to (3.4). Indeed, if 3.4 holds, then
@@
Next, let us describe the main idea behind our construction of a $r/2$-solution to (3.4). In what follows, the subcriticality of $L^{\infty}(0,T;\mathbb{L}^{r})$ with $r>1$ for the 2D Navier-Stokes equations (cf., the discussion below 1.2) plays a central role. Indeed, by subcriticality, given $q_{0}=\frac{r}{2}$, the solution $v_{0}$ to
@@
Let us now turn to the construction of a $q$-solution to (3.4). The above argument motivates the following splitting. Let $N$ be a positive integer such that
@@
At least formally, it is clear that $v$ solves (3.4). The latter fact is a straightforward consequence of the following identity (letting $v_{N}:=\overline{v}$ for simplicity)
$\sum_{i,j=0}^{N}B(v_{i},v_{j})$ $=\sum_{i=0}^{N}B(v_{i},v_{i})+\sum_{0\leq j<i\leq N}B(v_{i},v_{j})+\sum_{0\leq i<j\leq N}B(v_{i},v_{j})$
$=\sum_{i=0}^{N}B(v_{i},v_{i})+\sum_{i=0}^{N-1}\sum_{j=0}^{i-2}B(v_{i-1},v_{j})+\sum_{j=0}^{N-1}\sum_{i=0}^{j-2}B(v_{i},v_{j-1})$
$=B(\overline{v},\overline{v})+\sum_{i=0}^{N-1}\sum_{j=0}^{i-1}B(v_{i-1},v_{j})+\sum_{j=0}^{N-1}\sum_{i=0}^{j-2}B(v_{i},v_{j-1}).$
To show rigorously that $v$ given in (3.6) with $(v_{0},\dots,v_{N-1},\overline{v})$ solving (3.7) is a $q$-solution to (3.4) we need to check that $v_{0},\dots,v_{N-1}$ and $\overline{v}$ are sufficiently regular. The appropriate regularity class for $v$ in (3.6) to obtain a $q$-solution is given in the following definition, see also 3.6 below.
@@
We observe that $q_{i}>1$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. As an immediate consequence of 3.5, Sobolev embedding theorem and interpolation we have that
@@
In particular, $r_{i}>2$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. Therefore one can easily check that all the duality pairings in 3.5 are well defined. Moreover, for all $i\in\{1,\dots,N-1\}$,
@@
In particular, if $(v_{0},\dots,v_{N-1},\overline{v})$ is a $(p,r)$ solution of (3.7), then, given $v:=\sum_{i=0}^{N-1}v_{i}+\overline{v}$, $u=v+w$ is a $r/2$-solution of (1.2) in the sense of 1.2.
The following yields the well-posedness of (3.4) in the sense of 3.5.
@@
Let 3.4 be satisfied. For each $u_{\mathrm{in}}\in H$, $p\geq 2^{N}\frac{r}{r-2}$ there exists a unique $(p,r)$-solution $(v_{0},\dots,v_{N-1},\overline{v})$ of (3.7) in the sense of 3.5. Moreover $\overline{v}$ satisfies the energy relation
@@
If $\left(u_{\mathrm{in}}^{n}\right)_{n\in{\mathbb{N}}}$ is a sequence in $H$ converging to $u_{\mathrm{in}}\in H$ and $(w^{n})_{n\in{\mathbb{N}}}$ is a sequence in $C([0,T];\mathbb{L}^{r})$ converging to $w\in C([0,T];\mathbb{L}^{r})$, then the corresponding unique solutions $((v_{0}^{n},\dots,v_{N-1}^{n},\overline{v}^{n}))_{n\in{\mathbb{N}}}$ converge to the corresponding solution $(v_{0},\dots,v_{N-1},\overline{v})$, each one in the topologies of 3.5.
@@
Step 1: Linear part of (3.7). We argue by induction and exploit maximal $L^{p}$ regularity techniques, see [53, Chapter 3]. The existence and uniqueness of $v_{0}$ satisfying the corresponding PDE in the sense of 3.5 and the continuous dependence from data, i.e. $w$ in the topology of $C([0,T];\mathbb{L}^{r})$, follows if
@@
Now assume we have already proved the existence and uniqueness of $(v_{i})_{i\in\{0,\dots,l-1\}},$ $l\leq N-1$ solving (3.7) in the sense of 3.5 and depending continuously from the data, i.e. $w$ in the topology of $C([0,T];\mathbb{L}^{r})$. Let us check that there exists a unique $v_{l}$ solving the corresponding PDE in (3.7) in the sense of 3.5 and depending continuously from $w$ in the topology of $C([0,T];\mathbb{L}^{r})$. Again, due to maximal $L^{p}$ regularity techniques, it is enough to show that
@@
The claim is true, indeed due to 3.6
@@
By 2.6 and [29, Lemma 1.14], $\widetilde{f}\in L^{2}\left(0,T;V^{*}\right)$. Then, by Theorem 2.5,
@@
We conclude $z=0$ by the Gronwall lemma, using (3.15) and the integrability properties of $\overline{v}^{\left(2\right)}$.
@@
$\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}\left(b\left(\overline{v}^{n+1},\overline{\phi},\overline{v}^{n+1}\right)-b\left(\overline{v}^{n},\phi,\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s$
@@
$=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(z_{n},\tilde{v}\right)+B\left(\tilde{v},z_{n}\right)\right)\left(s\right),\phi\right\rangle\,\mathrm{d}s$
@@
As above, thanks to [29, Lemma 1.14] we deduce that
@@
Now we work under the bounds (3.17) and deduce, using the Gronwall lemma, for $\overline{T}$, depending only from $\lVert u_{\mathrm{in}}\rVert$ and $\lVert\tilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}$, possibly smaller than the previous one,
$\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n+1}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n+1}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s$
$\leq\frac{1}{2}\left(\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\right).$
It implies that the sequence $\left(\overline{v}^{n}\right)$ is Cauchy in $C\left(\left[0,\overline{T}\right];H\right)\cap L^{2}\left(0,\overline{T};V\right)$. The limit $\overline{v}$ has the right regularity to be a weak solution and satisfies the weak formulation; in the identity above for $\overline{v}^{n+1}$ and $\overline{v}^{n}$ we may prove that
$\int_{0}^{t}b\left(\overline{v}^{n+1}\left(s\right),\overline{\phi},\overline{v}^{n+1}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\overline{v}\left(s\right),\overline{\phi},\overline{v}\left(s\right)\right)\,\mathrm{d}s$
$\int_{0}^{t}b\left(\overline{v}^{n}\left(s\right),\overline{\phi},\tilde{v}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\overline{v}\left(s\right),\overline{\phi},\tilde{v}\left(s\right)\right)\,\mathrm{d}s$
$\int_{0}^{t}b\left(\tilde{v}\left(s\right),\overline{\phi},\overline{v}^{n}\left(s\right)\right)\,\mathrm{d}s$ $\rightarrow\int_{0}^{t}b\left(\tilde{v}\left(s\right),\overline{\phi},\overline{v}\left(s\right)\right)\,\mathrm{d}s.$
All these convergences can be proved easily by recalling the definition of $b$. Similarly, we can pass to the limit in the energy identity. After proving existence and uniqueness in $[0,\overline{T}]$ we can reiterate the existence procedure and in a finite number of steps cover the interval $[0,T]$.
@@
We can easily bound the right-hand side of relation (3.2) by Young’s inequality and Hölder’s inequality obtaining
@@
Applying Gronwall’s inequality to relation (3.2) the claimed continuity follows. ∎
@@
Combining 3.1, Theorem 3.7 and 3.8 we get immediately the existence of a $q-$solution of equation (1.2) in the sense of 1.2 for each $q\in(1,q_{\mathcal{H}})$.
@@
As discussed above, the results of subsection 3.1, subsection 3.2 provide the existence of a $q$-solution of equation (1.2) in the sense of 1.2 for each $q\in(1,q_{\mathcal{H}})$, moreover such a solution is adapted with paths in $C([0,T];H)$ due to 3.6. In order to conclude the proof of Theorem 1.31 it is enough to show the following:
@@
Let $u_{1}$ be a $q_{1}$-solution of (1.2) and $u_{2}$ be a $q_{2}$-solution of (1.2). Then $u_{1}\equiv u_{2}$.
The uniqueness result in the Ladyzhenskaya–Prodi–Serrin class of 3.9 might be known to experts. Here, for completeness, we provide a relatively short proof relying on maximal $L^{p}$-regularity techniques which seem not standard even in the absence of noise.
Proof of 3.9.
Recall that $u_{i}=w_{g}+v_{i}$ where $v_{i}$ solves the modified Navier-Stokes equations for $i\in\{1,2\}$ by 1.2. We split the proof into two cases.
@@
As $v_{i}\in L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ for $i\in\{1,2\}$ by 1.2, we obtain
@@
where in the last step we used the trace embedding [53, Theorem 3.4.8] applied with $A=A_{q}$.
@@
Hence, by maximal $L^{q}$-regularity of $A_{q}$, again the trace embedding [53, Theorem 3.4.8] as well as the Hölder inequality, there exists a constant $C_{0}>0$ independent of $v_{1},v_{2}$ and $\delta$ such that, for all $t\in[0,T]$ and $\mathbf{P}-a.s.$,
@@
Now the conclusion follows from the Grownall lemma and the integrability conditions on $v_{1},v_{2}$ due to 1.2 and $w_{g}\in C([0,\infty);L^{2q})$ for all $q<q_{\mathcal{H}}$ by 3.1.
@@
Let $w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$ and $2<\widetilde{r}\leq r$. If $(v_{0},\dots,v_{N-1},\overline{v})$ is a solution of (3.7) in the sense of 3.5 with
@@
and $(\widetilde{v}_{0},\dots,\widetilde{v}_{N-1},\widehat{v})$ is a solution of (3.4) in the sense of 3.5 with
@@
and we are left to show that $\overline{v}=\sum_{i=N}^{\widetilde{N}-1}\widetilde{v}_{i}+\widetilde{\overline{v}}=:\widehat{v}$. Due to previous steps we can assume that $v$ is $(\widetilde{p},r)$ solution since $\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2}>2^{{N}}\frac{{r}}{{r}-2}$ We observe that due to 3.5 and 3.6,
@@
Now the proof proceeds as in the first case of 3.9 and we omit the details. ∎

### v2

@@ 3.2. Auxiliary Navier-Stokes type equations
with $A_{q}$ is the Stokes operator on $\mathbb{L}^{q}$ and $B$ is the bilinear nonlinearity as in Section 2.3. Finally, $q=r/2$ and $(w,r)$ satisfies the following
@@
Note that the above assumption is satisfied $\mathbf{P}-a.s.$ with $w=w_{g}$, as it follows from Proposition 3.1. Moreover, the limitation $r<4$ is motivated by Remark 3.2. In particular, the arguments used in [1] do not apply to (3.10). Indeed, if 3.6 holds, then
@@
Next, let us describe the main idea behind our construction of a solution to (3.10). In what follows, the subcriticality of $L^{\infty}(0,T;\mathbb{L}^{r})$ with $r>1$ for the 2D Navier-Stokes equations (cf. the discussion below Definition 1.2) plays a central role. Indeed, by subcriticality, given $q_{0}=\frac{r}{2}$, the solution $v_{0}$ to
@@
Let us now turn to the construction of a solution to (3.10). The above argument motivates the following splitting. Let $N$ be a positive integer such that
@@
At least formally, it is clear that $v$ solves (3.10). The latter fact is a straightforward consequence of the following identity involving $B(v+w,v+w)$ (letting $v_{-1}=w,\ v_{N}:=\overline{v}$ for simplicity)
$B(v+w,v+w)$ $=\sum_{i,j=-1}^{N}B(v_{i},v_{j})$
$=\sum_{i=-1}^{N}B(v_{i},v_{i})+\sum_{i=-1}^{N}\sum_{j=-1}^{i-1}B(v_{i},v_{j})+\sum_{j=-1}^{N}\sum_{i=-1}^{j-1}B(v_{i},v_{j})$
${=B(\overline{v},\overline{v})+\sum_{i=-1}^{N-1}B(v_{i},v_{i})+\sum_{i=-1}^{N-1}B(\overline{v},v_{i})+\sum_{i=-1}^{N-1}\sum_{j=-1}^{i-1}B(v_{i},v_{j})}$
${+\sum_{i=-1}^{N-1}B(v_{i},\overline{v})+\sum_{j=-1}^{N-1}\sum_{i=-1}^{j-1}B(v_{i},v_{j})}$
$=B(\overline{v},\overline{v})+\sum_{i=-1}^{N-1}B(v_{i},\overline{v})+\sum_{i=-1}^{N-1}B(\overline{v},v_{i})$
$+\sum_{i=-1}^{N-1}\sum_{j=-1}^{i}B(v_{i},v_{j})+\sum_{j=-1}^{N-1}\sum_{i=-1}^{j-1}B(v_{i},v_{j})$
$=\left[B(\overline{v},\overline{v})+B\left(w+\sum_{j=0}^{N-1}v_{i},\overline{v}\right)+B\left(\overline{v},w+\sum_{j=0}^{N-1}v_{j}\right)\right.$
$\left.\quad\quad+B\left(v_{N-1},w+\sum_{j=0}^{N-1}v_{j}\right)+B\left(w+\sum_{j=0}^{N-2}v_{j},v_{N-1}\right)\right]$
$+\sum_{i=-1}^{N-2}\left[B\left(v_{i},w+\sum_{j=0}^{i}v_{j}\right)+B\left(w+\sum_{j=0}^{i-1}v_{j},v_{i}\right)\right].$
Looking at the last line, we can formally identify the first bracketed term as the one appearing in the equation for $\overline{v}$ and the $i$-th summand as the one appearing in the equation for $v_{i+1},\ i+1\in\{0,1,\dots,N-1\}.$ To show rigorously that $v$ given in (3.12) with $(v_{0},\dots,v_{N-1},\overline{v})$ solving (3.13) is a solution to (3.10) we need to check that $v_{0},\dots,v_{N-1}$ and $\overline{v}$ are sufficiently regular. The appropriate regularity class for $v$ in (3.12) to obtain a solution is given in the following definition, see also Remark 3.8 below.
@@
We observe that $q_{i}>1$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. As an immediate consequence of Definition 3.7, Sobolev embedding theorem and interpolation we have that
@@
In particular, $r_{i}>2$ for all $i\in\{0,\dots,N-1\}$ and is increasing in $i$. Therefore one can easily check that all the duality pairings in Definition 3.7 are well defined. Moreover, for all $i\in\{1,\dots,N-1\}$,
@@
In particular, if $(v_{0},\dots,v_{N-1},\overline{v})$ is a $(p,r)$ solution of (3.13), then, given $v:=\sum_{i=0}^{N-1}v_{i}+\overline{v}$, $u=v+w$ is a $r/2$-solution of (1.2) in the sense of Definition 1.2.
The following yields the well-posedness of (3.10) in the sense of Definition 3.7.
@@
Let 3.6 be satisfied. For each $u_{\mathrm{in}}\in H$, $p\geq 2^{N}\frac{r}{r-2}$ there exists a unique $(p,r)$-solution $(v_{0},\dots,v_{N-1},\overline{v})$ of (3.13) in the sense of Definition 3.7. Moreover $\overline{v}$ satisfies the energy relation
@@
If $\left(u_{\mathrm{in}}^{n}\right)_{n\in{\mathbb{N}}}$ is a sequence in $H$ converging to $u_{\mathrm{in}}\in H$ and $(w^{n})_{n\in{\mathbb{N}}}$ is a sequence in $C([0,T];\mathbb{L}^{r})$ converging to $w\in C([0,T];\mathbb{L}^{r})$, then the corresponding unique solutions $((v_{0}^{n},\dots,v_{N-1}^{n},\overline{v}^{n}))_{n\in{\mathbb{N}}}$ converge to the corresponding solution $(v_{0},\dots,v_{N-1},\overline{v})$, each one in the topologies of Definition 3.7.
@@
Step 1: Linear part of (3.13). We argue by induction and exploit maximal $L^{p}$ regularity techniques, see [55, Chapter 3]. The existence and uniqueness of $v_{0}$ satisfying the corresponding PDE in the sense of Definition 3.7 and the continuous dependence from data, i.e. $w$ in the topology of $C([0,T];\mathbb{L}^{r})$, follows if
@@
Now assume we have already proved the existence and uniqueness of $(v_{i})_{i\in\{0,\dots,l-1\}},$ $l\leq N-1$ solving (3.13) in the sense of Definition 3.7 and depending continuously from the data, i.e. $w$ in the topology of $C([0,T];\mathbb{L}^{r})$. Let us check that there exists a unique $v_{l}$ solving the corresponding PDE in (3.13) in the sense of Definition 3.7 and depending continuously from $w$ in the topology of $C([0,T];\mathbb{L}^{r})$. Again, due to maximal $L^{p}$ regularity techniques, it is enough to show that
@@
The claim is true, indeed due to Remark 3.8
@@
By Lemma 2.7, $\widetilde{f}\in L^{2}\left(0,T;V^{*}\right)$. Then, by Theorem 2.6,
@@
We conclude $z=0$ by the Grönwall lemma, using (3.21) and the integrability properties of $\overline{v}^{\left(2\right)}$.
@@
$\left\langle z_{n+1}\left(t\right),\overline{\phi}\right\rangle-\int_{0}^{t}\left(b\left(\overline{v}^{n+1},\overline{\phi},\overline{v}^{n+1}\right)-b\left(\overline{v}^{n},\overline{\phi},\overline{v}^{n}\right)\right)\left(s\right)\,\mathrm{d}s$
@@
$=-\int_{0}^{t}\left\langle z_{n+1}\left(s\right),A\overline{\phi}\right\rangle\,\mathrm{d}s-\int_{0}^{t}\left\langle\left(B\left(z_{n},\tilde{v}\right)+B\left(\tilde{v},z_{n}\right)\right)\left(s\right),\overline{\phi}\right\rangle\,\mathrm{d}s$
@@
As above, thanks to Lemma 2.7 we deduce that
@@
Now we work under the bounds (3.23) and deduce, using the Grönwall lemma, for $\overline{T}$, depending only from $\lVert u_{\mathrm{in}}\rVert$ and $\lVert\tilde{v}\rVert_{L^{\frac{2r}{r-2}}(0,T;\mathbb{L}^{r})}$, possibly smaller than the previous one,
$\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n+1}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n+1}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\leq\frac{1}{2}\left(\sup_{t\in\left[0,\overline{T}\right]}\lVert z_{n}\left(t\right)\rVert^{2}+\int_{0}^{\overline{T}}\lVert z_{n}\left(s\right)\rVert_{V}^{2}\,\mathrm{d}s\right).$
Now we can proceed as in the second step of the proof of [1, Theorem 3.3], showing that the sequence $\left(\overline{v}^{n}\right)$ is Cauchy in $C\left(\left[0,\overline{T}\right];H\right)\cap L^{2}\left(0,\overline{T};V\right)$. Its limit $\overline{v}$ is a weak solution of (3.22) on $[0,\overline{T}]$ and , hence, by the previous step, it is the unique solution. We refer the reader to [1, Theorem 3.3] for further details. After proving existence and uniqueness in $[0,\overline{T}]$ we can reiterate the existence procedure and in a finite number of steps cover the interval $[0,T]$.
@@
Thanks to 2.7, we can easily bound the right-hand side of relation (3.2) by Young’s inequality and Hölder’s inequality obtaining
@@
Applying Grönwall’s inequality to relation (3.2) the claimed continuity follows. ∎
@@
Combining Proposition 3.1, Lemma 3.5, Theorem 3.9 and Remark 3.10 we get immediately the existence of a $q$-solution of equation (1.2) in the sense of Definition 1.2 for each $q\in(1,q_{\mathcal{H}})$.
@@
As discussed above, the results of Section 3.1, Section 3.2 provide the existence of a $q$-solution of equation (1.2) in the sense of Definition 1.2 for each $q\in(1,q_{\mathcal{H}})$, moreover such a solution is adapted with paths in $C([0,T];H)$ due to Remark 3.8. Here we are left to discuss the problem of uniqueness. In order to reach our goal, we start providing a lemma which shows equivalence of $q$-solutions in the sense of 1.2 and those of the form $u=w_{g}+v$ as described by 3.11 below. Then we conclude the proof by providing uniqueness for solutions of the form $u=w_{g}+v$.
Lemma 3.11.
Let $T<\infty$, $u_{\mathrm{in}}\in L^{0}_{\mathscr{F}_{0}}(\Omega;\mathbb{L}^{2})$, $q\in(1,q_{\mathcal{H}})$ and assuming 1.1. Then $u$ is a $q$-solution to (1.2) in the sense of Definition 1.2 with paths in $C([0,T];\mathbb{L}^{2})$ $\mathbf{P}-a.s.$ if and only if $v:=u-w_{g}$ is progressively measurable with paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})\cap C([0,T];\mathbb{L}^{2})$ $\mathbf{P}-a.s.$ and it solves, for all divergence-free $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e. $t\in(0,T)$,
(3.26) $\int_{\mathcal{O}}v(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
Proof.
The proof is a trivial consequence of the notion of $q$-solution in 1.2, the regularity of $w_{g}$ in case of 1.1, i.e. 3.1, and the equivalence between weak and mild solutions for the linear stochastic problem, see 3.5. ∎
We call a $v$ progressively measurable with paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})\cap C([0,T];\mathbb{L}^{2})$ $\mathbf{P}-a.s.$ satisfying (3.26) in the sense of Lemma 3.11 an auxiliary $q$-solution of (3.10). The uniqueness of auxiliary $q$-solutions and their independence on $q$ is the content of the following result, which, combined with Lemma 3.11, concludes the proof of the first item in Theorem 1.3.
@@
Let $v_{1}$ be an auxiliary $q_{1}$-solution of (3.10) and $v_{2}$ be an auxiliary $q_{2}$-solution of (3.10). Then $v_{1}\equiv v_{2}$.
The uniqueness result in the Ladyzhenskaya–Prodi–Serrin class of Proposition 3.12 might be known to experts. Here, for completeness, we provide a relatively short proof relying on maximal $L^{p}$-regularity techniques which seem not standard even in the absence of noise.
Proof of Proposition 3.12.
We split the proof into two cases.
@@
As $v_{i}\in L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ for $i\in\{1,2\}$, we obtain
@@
where in the last step we used the trace embedding [55, Theorem 3.4.8] applied with $A=A_{q}$.
@@
Hence, by maximal $L^{q}$-regularity of $A_{q}$, again the trace embedding [55, Theorem 3.4.8] as well as the Hölder inequality, there exists a constant $C_{0}>0$ independent of $v_{1},v_{2}$ and $\delta$ such that, for all $t\in[0,T]$ and $\mathbf{P}-a.s.$,
@@
Now the conclusion follows from the Grönwall lemma and the integrability conditions on $v_{1},v_{2}$ and $w_{g}\in C([0,\infty);L^{2q})$ for all $q<q_{\mathcal{H}}$ by Proposition 3.1.
@@
Let $w\in C([0,T];\mathbb{L}^{r})$ for some $r\in(2,4)$ and $2<\widetilde{r}\leq r$. If $(v_{0},\dots,v_{N-1},\overline{v})$ is a solution of (3.13) in the sense of Definition 3.7 with
@@
and $(\widetilde{v}_{0},\dots,\widetilde{v}_{\widetilde{N}-1},\widetilde{\overline{v}})$ is a solution of (3.10) in the sense of Definition 3.7 with
@@
and we are left to show that $\overline{v}=\sum_{i=N}^{\widetilde{N}-1}\widetilde{v}_{i}+\widetilde{\overline{v}}=:\widehat{v}$. Due to previous steps we can assume that $v$ is $(\widetilde{p},r)$ solution since $\widetilde{p}\geq 2^{\widetilde{N}}\frac{\widetilde{r}}{\widetilde{r}-2}>2^{{N}}\frac{{r}}{{r}-2}$. We observe that due to Definition 3.7 and Remark 3.8,
@@
Now the proof proceeds as in the first case of Proposition 3.12 and we omit the details. ∎